### v2

@@ Methods
Similar to [44], low-order Finite Element techniques are used to discretize the Navier-Stokes Cahn-Hilliard equations. For the Navier-Stokes equations (3),(4) a pressure-correction projection method with second order backward differencing formula (BDF2) in time, described in [62], and a Taylor-Hood conforming Finite Element pair for in space are used. For the Cahn-Hilliard equation (5),(6) a conforming Finite Element pair for and Crank-Nicholson in time with mass lumping is used. The equations are decoupled by solving sequentially first the interface motion using the velocity from the previous time step followed by a Navier-Stokes solve using the updated phase variable. This works well for reasonably small time steps. The resulting linear systems are solved with a (S)SOR preconditioned GMRes or CG method from the DUNE-ISTL module [63] with a tolerance of . To reduce the computational effort, mesh adaptation based on the newest vertex bisection is applied [64, 65]. The refinement indicator is based on the phase variable and defined as . A mesh element is refined when and coarsened whenever . The implementation is done in the open-source framework DUNE and in particular DUNE-FEM [66, 67] using the Python based user interfaces, which allows the description of weak forms using UFL [68].
@@
As shown in Figure 1(b), 4D-ONIX comprises three networks. The first network is an encoder ( ), implemented using CNNs. The encoder sees all of the input projection images and extracts latent features of the sample from the projections. By transferring knowledge across different timestamps, it learns the general features of the sample, which is crucial for accurate 3D reconstruction from sparse projections. The second network is an IoR generator ( ), formed by a MLP. It learns the mapping from 4D spatial-temporal coordinates (x ,t ) to the refractive index of the sample ( ), with the assistance of the encoder. The generator takes as input the extracted latent features from the encoder and the 4D coordinates at each spatial-temporal point. It is trained to output the value of the refractive index at each 4D point. Predicted projections can be calculated by integrating the output refractive index along the line of propagation, following the law of X-ray propagation and interaction with matter [58]. The third network is a CNN discriminator ( ). The discriminator learns to minimize the difference between the image patches from the real projections and the predictions generated by 4D-ONIX. It sees both the image patches from the real projections and the 4D-ONIX predictions, and it learns to distinguish the fake ones from the real images. The encoder and the generator are optimized based on the feedback from the discriminator. They are trained to fool the discriminator by generating indistinguishable images, leading to high-quality 3D reconstruction of the sample.
@@
where and refer to image patches from the real and predicted projections, respectively. The simulation dataset included both absorption and phase contrast, resulting in two-channel representations of and . Since the phase contrast in our simulations was directly proportional to the absorption contrast, only the absorption images were shown in the present work. In the case of the experimental XMPI data, only absorption-contrast images were obtained. Therefore, and are representations of absorption contrast. The discriminator is trained to minimize the difference between these two patches. The data distribution, denoted as , represents the distribution over the collected projections in our experiments, which are considered real projections by the discriminator. On the other hand, represents the distribution over all generated predictions, which are considered fake projections by the discriminator. represents the expectation of a function. By minimizing the difference between the real and predicted projections, the discriminator provides feedback to the generator and the encoder networks, enabling them to generate more accurate and realistic 3D reconstructions.
@@
where and denote image patches from the real and predicted projections, respectively. and denote the two angles from which the projections were recorded, and stands for the view angle used in training. The adversarial loss was applied starting from the sixth epoch, once a convergence was obtained from the consistency of the two projections. The ADAM optimizer [72] with a mini-batch size of 8 was used throughout the training. We set the learning rates to be 0.0001 for all of the networks and decayed the learning rate by 0.1 every 100 epochs. The results presented in the 4D-ONIX performance evaluation section were the results of 200 epochs, which took around 70 hours of training on a NVIDIA V100 GPU with 32 GB of RAM.
@@
The data that support the findings of this study will be publicly available.
@@
The code developed in this study is available at https://github.com/yuhez/ONIX_XMPI_Recon.git.
@@ Acknowledgement
We are grateful to Z. Matej for his support and access to the GPU‐computing cluster at MAX IV. This work has received funding by ERC‐2020‐STG 3DX‐FLASH 948426 and the HorizonEIC-2021-PathfinderOpen-01-01, MHz-TOMOSCOPY 101046448.

### v3

@@ Methods
Similar to [50], low-order Finite Element techniques are used to discretize the Navier-Stokes Cahn-Hilliard equations. For the Navier-Stokes equations (3),(4) a pressure-correction projection method with second order backward differencing formula (BDF2) in time, described in [71], and a Taylor-Hood conforming Finite Element pair for in space are used. For the Cahn-Hilliard equation (5),(6) a conforming Finite Element pair for and Crank-Nicholson in time with mass lumping is used. The equations are decoupled by solving sequentially first the interface motion using the velocity from the previous time step followed by a Navier-Stokes solve using the updated phase variable. This works well for reasonably small time steps. The resulting linear systems are solved with a (S)SOR preconditioned GMRes or CG method from the DUNE-ISTL module [72] with a tolerance of . To reduce the computational effort, mesh adaptation based on the newest vertex bisection is applied [73, 74]. The refinement indicator is based on the phase variable and defined as . A mesh element is refined when and coarsened whenever . The implementation is done in the open-source framework DUNE and, in particular, DUNE-FEM [75, 76] using the Python based user interfaces, which allows the description of weak forms using UFL [77].
@@
As shown in Figure 1(b), 4D-ONIX comprises three networks. The first network is an encoder ( ), implemented using CNNs. The encoder sees all of the input projection images and extracts latent features of the sample from the projections. By transferring knowledge across different timestamps, it learns the general features of the sample, which is crucial for accurate 3D reconstruction from sparse projections. The second network is an IoR generator ( ), formed by a MLP. It learns the mapping from 4D spatial-temporal coordinates (x , t ) to the refractive index of the sample (, ) , with the assistance of the encoder. Here, and are real and positive numbers representing the real and imaginary parts of the complex refractive index, respectively. The generator takes as input the extracted latent features from the encoder and the 4D coordinates at each spatial-temporal point. It is trained to output the value of the refractive index at each 4D point. Predicted projections can be calculated by integrating the output refractive index along the line of propagation, following the law of X-ray propagation and interaction with matter [64]. Figure 7 illustrates the process of novel view prediction. The third network is a CNN discriminator ( ). The discriminator learns to minimize the difference between the image patches from the real projections and the predictions generated by 4D-ONIX. It sees both the image patches from the real projections and the 4D-ONIX predictions, and it learns to distinguish the fake ones from the real images. The encoder and the generator are optimized based on the feedback from the discriminator. They are trained to fool the discriminator by generating indistinguishable images, leading to high-quality 3D reconstruction of the sample.
@@
where and refer to image patches from the real and predicted projections, respectively. The simulation dataset included both absorption and phase contrast, resulting in two-channel representations of and . Since the phase contrast in our simulations was directly proportional to the absorption contrast, only the absorption images were shown in the present work. In the case of the experimental data, only absorption-contrast images were obtained. Therefore, and are representations of absorption contrast. The discriminator is trained to minimize the difference between these two patches. The data distribution, denoted as , represents the distribution over the collected projections in our experiments, which are considered real projections by the discriminator. On the other hand, represents the distribution over all generated predictions, which are considered fake projections by the discriminator. represents the expectation of a function. By minimizing the difference between the real and predicted projections, the discriminator provides feedback to the generator and the encoder networks, enabling them to generate more accurate and realistic 3D reconstructions.
@@
where and denote image patches from the real and predicted projections, while and stand for the view angle of the recorded projections and the predictions, respectively. The adversarial loss was applied starting from the sixth epoch, once a convergence was obtained from the consistency of the constraining projections. The ADAM optimizer [81] with a mini-batch size of 8 was used throughout the training. For the training of water droplet datasets, We set the learning rates to be 0.0001 for the networks and decayed the learning rate by 0.1 every 100 epochs. The results presented in the 4D-ONIX demonstration on simulated water droplet collisions section were the results of 200 epochs, which took around 70 hours of training on a NVIDIA V100 GPU with 32 GB of RAM. For the experimental additive manufacturing data, we employed a random dice mode for training, where half of the iterations used GAN loss and the other half used MSE loss. The learning rate was set to 0.0001, and no decay was applied. The results presented in the 4D-ONIX demonstration on experimental additive manufacturing data section were based on 400 epochs of training, which took 35 hours on an NVIDIA A100 GPU with 80 GB of RAM.
@@
The data that support the findings of this study is available at figshare.
@@
The code developed in this study is available at https://github.com/yuhez/4D-ONIX.
@@ Acknowledgement
We are grateful to Z. Matej for his support and access to the GPU‐computing cluster at MAX IV. We are also grateful to Malgorzata Makowska, et al., for sharing the additive manufacturing data [68] to demonstrate the capabilities of 4D-XMPI. This data was acquired within a frame of the project “Operando tomography of Selective Laser Additive Manufacturing” funded from the SNF Spark grant CRSK-2_196085. This work has received funding by ERC‐2020‐STG 3DX‐FLASH 948426 and the HorizonEIC-2021-PathfinderOpen-01-01, MHz-TOMOSCOPY 101046448.